### v1

@@ 8 Pulsar-based positioning and time-keeping
Acknowledgements. We thank Brynmor Haskell for providing helpful comments. PZ acknowledges the support from NSFC grant No. 12273010. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). JHF is supported by the China National Natural Science Foundation (No. 12433004). YC acknowledges support from a Ramon y Cajal fellowship (RYC2021-032718-I) financed by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. BDM acknowledges support via a Ramón y Cajal Fellowship RYC2018-025950-I and the Spanish MINECO grant PID2022-136828NB-C44. BDM, YC and GS acknowledge support from the Spanish MINECO (PID2023-148661NB-I00), the E.U. FEDER funds, and the AGAUR/Generalitat de Catalunya (SGR-386/2021). LD acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 549824807. PR acknowledges support by the “Programma di Ricerca Fondamentale INAF 2023”. YH is supported by National Natural Science Foundation of China (grant No. 12233002), by National SKA Program of China No. 2020SKA0120300, by the National Key R&D Program of China (2021YFA0718500), and by the support from the Xinjiang Tianchi Program. C-YN is supported by a GRF grant of the Hong Kong Government under HKU 17304524. FX is supported by National Natural Science Foundation of China (grant No. 12373041 and No. 12422306), and Bagui Scholars Program (XF).

### v2

@@ 8 Pulsar-based positioning and time-keeping
Conflict of interests. The authors declare that they have no conflict of interest.
Acknowledgements. We thank B. Haskell, J.E. Horvath, N. Dutra Pire, B.B. Martins, M.G.B. de Avellar, L.M.de Sá, D. Vasconcelos Rodrigues for providing helpful comments. This work is supported by China’s Space Origins Exploration Program. PZ acknowledges the support from the National Natural Science Foundation of China (No. 12273010). S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). JHF is supported by the China National Natural Science Foundation (No. 12433004). AP acknowledges support from grant PID2021-124581OB-I0, PID2024-155316NB-I00 and 2021SGR00426. YC acknowledges support from a Ramon y Cajal fellowship (RYC2021-032718-I) financed by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. BDM acknowledges support via a Ramón y Cajal Fellowship RYC2018-025950-I and the Spanish MINECO grant PID2022-136828NB-C44. BDM, YC and GS acknowledge support from the Spanish MINECO (PID2023-148661NB-I00), the E.U. FEDER funds, and the AGAUR/Generalitat de Catalunya (SGR-386/2021). LD acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 549824807. PR acknowledges support by the “Programma di Ricerca Fondamentale INAF 2023”. YH is supported by National Natural Science Foundation of China (grant No. 12233002), by National SKA Program of China No. 2020SKA0120300, by the National Key R&D Program of China (2021YFA0718500), and by the support from the Xinjiang Tianchi Program. C-YN is supported by a GRF grant of the Hong Kong Government under HKU 17304524. FX is supported by National Natural Science Foundation of China (grant No. 12373041 and No. 12422306), and Bagui Scholars Program (XF).